Find all the real zeros of function f(x)=12x^4+35x^3+144x^2+114x+20
I already have x= -2/3

Answer by stanbon(75887)

Find all the real zeros of function f(x)=12x^4+35x^3+144x^2+114x+20
I already have x= -2/3
Also x=-1/4 is a zero
----------
Therefore (4x+1)(3x+2) =12x^2+11x+2 divides f(x)
Using long division you can find the remaining quotient, which is:
Quotient: x^2+2x+10
Use the quadratic formula to find the last two zeroes:
x=[-2+-sqrt(4-4(10)]/2
x=[-2+-6i]/2
x=-1+3i or x=-1-3i
These are the missing zeroes.
Cheers,
Stan H.